solve the pairsof equations using the addition or subtraction method:
2x-y=4 x+y=2
x-2y=6 3x+4y=8

Answer by checkley77(12844)

2x-y=4
x+y=2 now add
--------------------------
3x=6
x=6/3
x=2 answer.
2+y=2
y=2-2
y=0 answer.
----------------------
x-2y=6 multiply this equation by 2 & add.
3x+4y=8
2x-4y=12
-----------------
5x=20
x=20/5
x=4 answer.
4-2y=6
-2y=6-4
-2y=2
y=2/-2
y=-1 answer.